OpenStudy (mrlightbody):

find the value of the integral 1_^64 ((log4 x)^2 −√3 log4 x +7)/x dx

OpenStudy (anonymous):

put \(u=\log_4(x)\)

OpenStudy (anonymous):

change the limit of integration while you are at it, so you don't have to change back at the end \\[u=\log_4(x),u(1)=\log_4(1), u(64)=\log_4(64)\] both are cooked up to be easy

OpenStudy (mrlightbody):

⌠64 ⌡1 ((log4 x)^2 −√3 log4 x +7)/x dx i am still abit lost @satellite

OpenStudy (anonymous):

you know the derivative of \(\log_4(x)\)?

OpenStudy (anonymous):

if you do not, i can tell you

OpenStudy (mrlightbody):

no i dont

OpenStudy (anonymous):

do you know the derivative of \(\ln(x)\)?

OpenStudy (mrlightbody):

1/x right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and by the all mighty "change 'o base" formula \[\log_4(x)=\frac{\ln(x)}{\ln(4)}\] so the derivative is \[\frac{1}{x\ln(4)}\]

OpenStudy (anonymous):

keeping in mind that \(\ln(4)\) is just a constant (some number)

OpenStudy (anonymous):

make the \(u\) sub and you have it in one step basically

OpenStudy (anonymous):

\[\frac{1}{\ln(4)}\int_0^3(u^2-\sqrt{3}u+7)du\] in other words

OpenStudy (mrlightbody):

I understand now, then from taking out that constant I will be able to anti differentiate?

OpenStudy (mrlightbody):

@satellite

OpenStudy (anonymous):

yes use the backwards power rule for each term

OpenStudy (anonymous):

i get the feeling that the process was a bit of a mystery want to do it real slow?

OpenStudy (mrlightbody):

yeah that will be great, would make me understand even more

OpenStudy (mrlightbody):

what did you do with the x the divides the term ?

OpenStudy (anonymous):

it is a set up for a u-sub because each term has a \(\log_4(x)\) in it plus the derivative of \(\log_4(x)\) is right there in front of you, although hiding a little bit

OpenStudy (anonymous):

\[\int\frac{\log_4(x)^2-\sqrt3\log_4(x)+7}{x}dx\]

OpenStudy (anonymous):

think of it like this \[\int \log_4(x)-\sqrt3\log_4(x)-7)\times \frac{dx}{x}\]

OpenStudy (anonymous):

well actually like this \[\int( \log_4(x)^2-\sqrt3\log_4(x)-7)\times \frac{dx}{x}\]

OpenStudy (anonymous):

\(\frac{1}{x}\) is almost the derivative of \(\log_4(x)\) put \(u=\log_4(x)\) then \[du=\frac{dx}{x\ln(4)}\]

OpenStudy (anonymous):

the constant in the denominator doesn't make any difference, it comes right out front you get \[\frac{1}{\ln(4)}\int (u^2-\sqrt{3}u+7)du\]

OpenStudy (anonymous):

i am assuming that you have seen a u-sub before, just not one as not so obvious as this one for example if you had \[\int \sin^2(x)\cos(x)dx\] y ou would make \[u=\sin(x),du=\cos(x)dx\] and then integrate \[\int u^2du\]

OpenStudy (mrlightbody):

understand that part now

OpenStudy (anonymous):

then while you are at it, change the limits of integration so you don't have to revert back to \(\log_4(x)\) at the end you got that part?

OpenStudy (mrlightbody):

what do you mean change the limits ?

OpenStudy (mrlightbody):

@satellite

OpenStudy (mrlightbody):

thanks i clearly understand it and got the ans @satellite